-11-(x+30)=-41-2(5x-18). How do I solve this. Could someone show me the steps so that I can understand this problem, thanks.
Answer by algebrahouse.com(1659)

-11 - (x + 30) = -41 - 2(5x - 18)
-11 - x - 30 = -41 - 10x + 36 {used distributive property}
-x - 41 = -10x - 5 {combined like terms}
9x - 41 = -5 {added 10x to both sides}
9x = 36 {added 41 to both sides}
x = 4 {divided both sides by 9}
www.algebrahouse.com